OpenStudy (gabylovesyou):

Deon is putting money into a checking account. Let y represent the total amount of money in the account (in dollars). Let x represent the number of weeks Deon has been adding money. Suppose that x and y are related by the equation y =550+50x What is the change per week in the amount of money in the account? What was the starting amount of money in that account ?

OpenStudy (gabylovesyou):

@mathstudent55

OpenStudy (sshayer):

from 50 x you can imagine how much is added per week.

OpenStudy (sshayer):

if x=0 what is the amount?

OpenStudy (sshayer):

if x=1 50 x=?

OpenStudy (gabylovesyou):

50?

OpenStudy (sshayer):

that means 50 dollars are added per week

OpenStudy (sshayer):

wen x=0,y=?

OpenStudy (sshayer):

correction when x=0,y=?

OpenStudy (gabylovesyou):

ok so whats the starting number

OpenStudy (sshayer):

y=550+50x

OpenStudy (sshayer):

when x=0 y=?

OpenStudy (sshayer):

This gives you initial amount

OpenStudy (mathstudent55):

y = 550 + 50x No matter what x is, there is always a 550 in the equation. The amount $550 does not change. If x = 0 (at the beginning), or if x = 1 (at the end of the 1st week), or for any other x value, there is always 550 there. That means the account already had $550 before he started adding $50 per week. The 50 is being multiplied by x. If x = 0, then 50x = 50(0) = 0. At week 0, he has added 0 to the $550. At week 1, x = 1, and 50x = 50(1) = 50. At week 1 he has added $50. At week 2, x = 2, and 50x = 50(2) = 100. At week 2 he has added $100. $50 from week 1 and $50 more from week 2. At week 3, 50x = 50(3) = 150, so at week 3 he has added $150 from the 3 weekly $50 deposits. This should make clear how to answer the two questions.

OpenStudy (sshayer):

after 1 week amount=550+50*1=600 after 2 week amount=550+50*2=550+100=650 after 3 week amount 550+50*3=550+150=700 ....... 700-650=650-600=50 so amount added each week=50 dollar.